OpenStudy (slickster256):

how would i find the area of a circle with a diameter of 8

OpenStudy (anonymous):

Pi4^2

OpenStudy (xmissalycatx):

You use the equation Area = \[\pi * r ^{2}\]

OpenStudy (slickster256):

wouldnt it be (pi/4)X(diameter)^2

OpenStudy (texaschic101):

area of a circle = pi * r^2 r = 1/2 of d.....so if d is 8, then r = 4 A = pi * 4^2 A = pi * 16

OpenStudy (xmissalycatx):

So the radius is always half of the diameter, therefore, Area would be pi times 16 because 4 squared is 16!

OpenStudy (slickster256):

oh ok